A simple Harmonic oscillator has time period The the time taken by it to travel from the extreme position to half the amplitude is

Grade:

1 Answers

Arun
5 years ago
Dear student
 
As we know
x = a cos(wt)
Now
a/2 = a cos(wt)
cos (wt) = cos (pi/3)
wt = pi/3
2*pi* t /T  = pi/3
t = T/6
 
Regards
Arun (askIITians forum expert)